Jay buys a 30-year annuity immediate with end of year payments of 8,000 for a price of P. He replaces his capital over 30 years with a savings account, which pays AEIR 6%. His APY is 5%. Please find P.
Answer #1

Suppose the price of the annuity is equal to "P"

APY=((Payments/AEIR*[(1+AEIR)^n-1)]/P)^(1/n)-1
Here, n is the time period.
Prefers to the price of the annuity.

5%=((8000/6%*(1.06^30-1))/P)^(1/30)-1
5%=((133333.3333*(4.743491173))/P)^(1/30)-1
5%=(632465.4896/P)^(1/30)-1
1.05=(632465.4896/P)^(1/30)
(1.05)^30=(632465.4896/P)
4.321942375=(632465.4896/P)
P=632465.4896/4.321942375
So, the value of P=146338.2514


answered by: BugLittle